y=x^2-9x find the x intercepts
Found 2 solutions by sofiyac, nerdybill:
Answer by sofiyac(983)

set y=0 and solve for x
0=x(x-9)
x=0
x-9=0
x=9
So x intercepts are x=0 and x=9

Answer by nerdybill(7384)

y=x^2-9x find the x intercepts
.
Set y to zero and solve for x:
0=x^2-9x
0=x(x-9)
.
x = {0,9}
.
So, the x-intercepts are at:
(0,0) and (9,0)